# Eulerian-type polynomials over matchings and matching permutations

### v1

@@ 1. Introduction
The Catalan number $C_{n}=\frac{1}{n+1}\binom{2n}{n}$ counts matchings in $\mathcal{M}_{n}$ with no crossings (also known as noncrossing matchings), see [55, Exercise 6.19]. The Narayana number $N(n,k)=\frac{1}{n}\binom{n}{k-1}\binom{n}{k}$ counts noncrossing matchings in $\mathcal{M}_{n}$ with $k$ blocks of the form $(i,i+1)$, see [12, p. 3] for instance. It is well known that Eulerian numbers possess many of the same or similar properties as Narayana numbers, see [24, 41, 48] for instance. It is natural to investigate the relationship between Eulerian numbers and matchings.
@@
In Section 3, we provide a symmetric expansion of a five-variable neighbor polynomial of matchings. We also establish the relationship between neighbor polynomials and trivariate second-order Eulerian polynomials. In particular, by Corollary 21, we can get the following result.
@@ 2.1. Definitions and main results
When $y=1$, the polynomial $A_{n}(x,y)$ reduces to the classical Eulerian polynomial $A_{n}(x)$, which first appeared in a summation formula due to Euler:
@@
As observed in [6, 42, 43, 52, 54], the polynomials $A_{n}(x,p,q)$ contain a great deal of information about permutations and signed permutations. For example, the $r$-colored Eulerian polynomials introduced by Steingrímsson [56] can be expressed as follows (see [42, p. 30]):
@@
A few years later, Ren [49] investigated seven sets with the cardinality $a_{n}$, including the set of signed permutations with decreasing blocks all of whose left-to-right-minima have positive signs. Recently, Sokal and Zeng [54, Section 4] obtained $S$-fractions for perfect matchings by involving right-to-left minima and left-to-right maxima. They also observed some connections between matchings and permutations, see [54, Eq. (4.5)] for instance. As a continuation of these work, we introduce further definitions. For $M\in\mathcal{M}_{n}$, a block $\alpha=(i,j)$ is called
@@
The $p=x$ and $q=1$ case of Corollary 4 was also discussed in [40]. In [9, Corollary 7.4], Brenti discovered that $A_{n}(x,1,-1)=-(x-1)^{n-1}$. Subsequently, Kasraoui and Zeng [36] provided bijective proofs for $A_{n}(x,1,-1)=-(x-1)^{n-1}$ and $A_{n}(x,0,-1)=-x-x^{2}-\cdots-x^{n-1}$. Therefore, the following result is immediate.
@@ 2.3. $q$-derangement polynomials and $(p,q)$-Eulerian polynomials of type $B$
Let ${\rm cda\,}(\pi):=\#\{i:\pi^{-1}(i)<i<\pi(i)\}$ be the number of cycle double ascents of $\pi$. Let $\mathcal{D}_{n,k}:=\{\pi\in\mathcal{D}_{n}:{\rm cda\,}(\pi)=0,~{\rm exc\,}(\pi)=k\}$. Using the theory of continued fractions, Shin and Zeng [52, Theorem 11] obtained that
@@
More generally, combining (5) and [42, Theorems 3.4], we obtain the following result.
@@ 3.1. Definitions and preliminaries
    Sokal and Zeng [54, Section 4.2] studied crossings and nestings according to whether the element in second position is even or odd, i.e., a crossing or nesting $i<j<k<\ell$ is even (resp. odd) if $j$ is even (resp. odd).
@@ 3.2. Main results
It is well known that the coefficients $\alpha_{n,i}$ also counts 0-1-2 increasing plane trees on $[n]$ with $i$ degree two vertices, see [21] and [53, A101280] for instance. Clearly, a 0-1-2-3 increasing plane tree reduces to a 0-1-2 increasing plane tree if there are no degree three vertices, see Figure 3 for instance. It follows from Theorem 18 that
@@ 3.3. Relationship to trivariate second-order Eulerian polynomials
which was proved by Levande [38, Appendix A] using an inductive proof. In [12], Cameron and Killpatrick discovered that

### v2

@@ 1. Introduction
The Catalan number $C_{n}=\frac{1}{n+1}\binom{2n}{n}$ counts matchings in $\mathcal{M}_{n}$ with no crossings (also known as noncrossing matchings), see [56, Exercise 6.19]. The Narayana number $N(n,k)=\frac{1}{n}\binom{n}{k-1}\binom{n}{k}$ counts noncrossing matchings in $\mathcal{M}_{n}$ with $k$ blocks of the form $(i,i+1)$, see [12, p. 3] for instance. It is well known that Eulerian numbers possess many of the same or similar properties as Narayana numbers, see [24, 42, 49] for instance. It is natural to investigate the relationship between Eulerian numbers and matchings.
@@
In Section 3, we provide a symmetric expansion of a five-variable neighbor polynomial of matchings. We also establish the relationship between the neighbor polynomials and the trivariate second-order Eulerian polynomials. In particular, by Corollary 21, we get the following result.
@@ 2.1. Definitions and main results
When $y=1$, the polynomial $A_{n}(x,y)$ reduces to the Eulerian polynomial $A_{n}(x)$, which first appeared in a summation formula due to Euler:
@@
As observed in [6, 38, 43, 44, 53, 55], the polynomials $A_{n}(x,p,q)$ contain a great deal of information about permutations and signed permutations. For example, the $r$-colored Eulerian polynomials introduced by Steingrímsson [57] can be expressed as follows (see [43, p. 30]):
@@
A few years later, Ren [50] investigated seven sets with the cardinality $a_{n}$, including the set of signed permutations with decreasing blocks all of whose left-to-right-minima have positive signs. Recently, Sokal and Zeng [55, Section 4] obtained $S$-fractions for perfect matchings by involving right-to-left minima and left-to-right maxima. They also observed some connections between matchings and permutations, see [55, Eq. (4.5)] for instance. As a continuation of these work, we introduce further definitions. For $M\in\mathcal{M}_{n}$, a block $\alpha=(i,j)$ is called
@@
The $p=x$ and $q=1$ case of Corollary 4 was also discussed in [41]. In [9, Corollary 7.4], Brenti discovered that $A_{n}(x,1,-1)=-(x-1)^{n-1}$. Subsequently, Ksavrelof and Zeng [38] provided bijective proofs for $A_{n}(x,1,-1)=-(x-1)^{n-1}$ and $A_{n}(x,0,-1)=-x-x^{2}-\cdots-x^{n-1}$. Therefore, the following result is immediate.
@@ 2.3. $q$-derangement polynomials and $(p,q)$-Eulerian polynomials of type $B$
Let ${\rm cda\,}(\pi):=\#\{i:\pi^{-1}(i)<i<\pi(i)\}$ be the number of cycle double ascents of $\pi$. Let $\mathcal{D}_{n,k}:=\{\pi\in\mathcal{D}_{n}:{\rm cda\,}(\pi)=0,~{\rm exc\,}(\pi)=k\}$. Using the theory of continued fractions, Shin and Zeng [53, Theorem 11] obtained that
@@
More generally, combining (5) and [43, Theorems 3.4], we obtain the following result.
@@ 3.1. Definitions and preliminaries
    Sokal and Zeng [55, Section 4.2] studied crossings and nestings according to whether the element in second position is even or odd, i.e., a crossing or nesting $i<j<k<\ell$ is even (resp. odd) if $j$ is even (resp. odd).
@@ 3.2. Main results
It is well known that the coefficients $\alpha_{n,i}$ also counts 0-1-2 increasing plane trees on $[n]$ with $i$ degree two vertices, see [21] and [54, A101280] for instance. Clearly, a 0-1-2-3 increasing plane tree reduces to a 0-1-2 increasing plane tree if there are no degree three vertices, see Figure 3 for instance. It follows from Theorem 18 that
@@ 3.3. Relationship to trivariate second-order Eulerian polynomials
which was proved by Levande [39, Appendix A] using an inductive proof. In [12], Cameron and Killpatrick discovered that